Homework Assignment 1 You must show all your work to earn points ECON 3125 SP19 Name: 2. Suppose that a firm faces the demand curve, P- 300 - 40, where P denotes price in dollars and Q denotes total unit sales. The cost equation is TC-300+920 a. Determine the firms profit-maximizing output and price. 2 points b. Given the output (Q) value from part a, compute Total cost and Marginal cost when the cost equation is TC- 300+92Q: I point c. Suppose that there is a change in the production process so that the cost equation becomes TC = 1.800 + 10Q + Q, Compute the new Total cost and Marginal cost at the same output level (Q value) from part a: 1 point Determine the firms new output and price level when the cost equation is TC = 1,500 + 100+Q and the demand curve remains P 300-40 d. 2 points e. Using your answers to part b and c, explain why the profit maximizing quantity changed in I point your answers to part a and part d: Due date: Thursday 1/31 at the time of class
Homework Assignment 1 You must show all your work to earn points ECON 3125 SP19 Name: 3. A U.S. textbook publisher is introducing a new economics textbook, Managerial Economics I is no Graphing matter, to the domestic market. Each book is produced at a constant marginal cost of 598 per book. Management predicts that annual domestic demand for the book is P.- 278-0.3Q, where Po- price of a book in dollars, and Q denotes the number of books (as measured in thousands). Assuming no costs beyond the MC of $98 per book, state the profit equation for the domestic market: a. 1 polet b. State the marginal profit equation for the domestic market: 1 point c. Determine the optimal quantity and price of the book in the domestic market: 2 points d. Using your answers to parts a and b, compute the total profit in the domestic market assuming no fixed costs: I point 4. No Name University has a successful football team, the Lethargic Leeches, and sells tickets to students, alumni, and the public. Experience has shown that attendance has followed the demand relationship: Q-28,000-2,000P 1 point IfNo Name U charges $7 per ticket, predict attendance (Q). a. c. Now, assume that the current capacity of the football stadium is 10,000 seats, what tieket I point price should be set for the stadium to be sold-out? Due date: Thursday 1/31 at the time of class
Please show work!
Answer #1

1.

From the demand curve we can say that,

When Q=0, P=80 (vertical intercept) and when P=0, Q=8(Horizontal intercept)

a) Px = a - bQx

By using graph data,

80 = a - b x 0

a=80

and similarly,

0 = a - 8b

a=8b

then,

80 = 8b

b = 10

The equation for demand curve is: px = 80 - 10Qx

b) px = 80 - 10Qx

10Qx = 80 - Px

Qx = 8 - 0.1Px

c) Here total revenue (TR) = (Px).(Qx)

TRx = Qx.(80 - 10Qx)

The equation for total revenue is:

TRx = 80Qx - 10Qx2

d) The marginal revenue (MR) = dTR/dQx

MR = d/dQx(80Qx - 10Qx2) = 80 - 20Qx.

This is the equation for marginal revenue.